# Self-Supervised Multi-Part Articulated Objects Modeling via Deformable Gaussian Splatting and Progressive Primitive Segmentation

### v1

@@ 3.3 Coarse Part-Level Segmentation of Gaussian Primitives
Motion-driven separation of static and dynamic Gaussians.
Without object masks or semantic priors, we detect dynamic primitives by computing the maximum Euclidean displacement of each Gaussian center across the $K$ interaction states. Let $\boldsymbol{\mu}_{i}^{(k)}\in\mathbb{R}^{3}$ be the center of Gaussian primitives $g_{i}$ in state $k$. We define its maximal displacement is
$d_{i}=\max_{j,k\in\{1,\dots,K\}}\bigl{\lVert}\boldsymbol{\mu}_{i}^{(j)}-\boldsymbol{\mu}_{i}^{(k)}\bigr{\rVert}_{2},\quad\hat{d}_{i}=d_{i}/\!\max_{r}d_{r}\in[0,1].$ (8)
Gaussians with $\hat{d}_{i}\geq\tau_{\text{mot}}$ form the dynamic set $\mathcal{G}_{\text{dyn}}$, and the remainder $\mathcal{G}_{\text{stat}}$ is static, where the threshold $\tau_{\text{mot}}\in[0,1]$ specifies the fraction of the maximum displacement of the entire scene that qualifies as motion.
Estimating the number of motion parts via a VLM.
To infer how many rigid parts appear in $\mathcal{G}_{\text{dyn}}$ we sample $M$ image pairs $(I_{k_{i}}^{(v_{p})},I_{k_{j}}^{(v_{q})})$ with different states $k_{i}\neq k_{j}$. A visual language model (VLM) [39] is queried with the fixed prompt:
@@
Trajectory-based clustering of dynamic Gaussians.
For each $g_{i}\!\in\!\mathcal{G}_{\text{dyn}}$ we assemble a motion descriptor
$\mathbf{f}_{i}=\bigl{[}\hat{\Delta}\boldsymbol{\mu}_{i}^{(1)},\,\lVert\Delta\boldsymbol{\mu}_{i}^{(1)}\rVert_{2},\,\dots,\,\hat{\Delta}\boldsymbol{\mu}_{i}^{(K-1)},\,\lVert\Delta\boldsymbol{\mu}_{i}^{(K-1)}\rVert_{2}\bigr{]}\in\mathbb{R}^{4(K-1)},$ (9)
@@
$\Delta\boldsymbol{\mu}_{i}^{(k)}=\boldsymbol{\mu}_{i}^{(k+1)}-\boldsymbol{\mu}_{i}^{(k)},\qquad\hat{\Delta}\boldsymbol{\mu}_{i}^{(k)}=\frac{\Delta\boldsymbol{\mu}_{i}^{(k)}}{\lVert\Delta\boldsymbol{\mu}_{i}^{(k)}\rVert_{2}+\varepsilon}.$ (10)
Unit directions encode orientation and lengths keep relative travel, so the descriptor is translation/rotation aware yet scale balanced. After normalization of $\ell_{2}$, we cluster the descriptors with K-means [40] using $K=n_{\text{parts}}$. On the unit sphere, Euclidean distance equals cosine distance, therefore Gaussians that share one rigid motion, even at different amplitudes, are grouped together.
Clusters containing fewer than $2\%$ of dynamic Gaussians are regarded noise and merged into the nearest larger cluster. Each dynamic primitive then receives a part label $c_{i}\in\{1,\dots,n_{\text{parts}}\}$, while static ones are labeled 00. These coarse labels initialize the fine-grained stage that follows.
@@ 3.4 Refined Part-Level Segmentation of Gaussian Primitives
Visibility-aware prompt generation for SAM.
To sharpen the coarse part labels in Sec. 3.3 we first render each part through a visibility filter and then draw sparse point prompts for the Segment Anything Model (SAM) [32].
Per-Gaussian contribution at a pixel. For a pixel $\mathbf{x}$ in view $v$, the contribution weight $w_{i}^{(v)}(\mathbf{x})$ from Gaussian $g_{i}=(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$ is obtained by first evaluating its screen-space density $\rho_{i}^{(v)}(\mathbf{x})$ (Eq. (2)) and then applying the front-to-back compositing rule (Eq. (3)), which together map every three-dimensional primitive to a continuous weight field whose value at $\mathbf{x}$ equals $w_{i}^{(v)}(\mathbf{x})$.
@@
where $\tau_{\text{vis}}=1.2$ and $\varepsilon$ avoids division by zero. The winner margin test keeps only pixels in which part $p$ clearly dominates, thus reducing the ambiguity at overlaps.
@@
Boundary-aware Gaussian splitting.
For each Gaussian $g_{i}$ we find its nearest view $v_{\text{near}}$ by ray sampling and project the ellipsoid to a two-dimensional ellipse. If the ellipse major axis extends beyond the mask of the part $\mathcal{M}_{p}^{(v_{\text{near}})}$, the Gaussian becomes a boundary candidate.
We follow SAGD [36] to split each candidate, as shown in the bottom part of Fig. 1. The fraction of the major axis that lies inside the mask is measured in image space and the same ratio is applied to the center $\boldsymbol{\mu}_{i}$ and the scale vector $\mathbf{s}_{i}$, resulting in two children $g_{i}^{(\text{part})}$ and $g_{i}^{(\text{bg})}$. The part child retains the in-mask portion and the background child captures the overflow. The background child $g_{i}^{(\text{bg})}$ is re-evaluated in its nearest view; if its ellipse still crosses another part mask the split is applied recursively until every descendant lies inside exactly one mask.
Part-aware Gaussian field via local refinement.
Finally, parameters of unsplit Gaussians remain fixed, whereas all new children are locally fine-tuned for a few iterations on $\mathcal{L}_{\text{total}}$ (updating $\mathbf{q}$ and $\alpha$ only) to restore the photo consistency. The recursive split-and-tune process produces sharply aligned, mask-consistent part boundaries while leaving the converged Gaussian field undisturbed.
The entire procedure yields a refined, part-aware Gaussian field with inteactive state
$\mathcal{G}^{(\boldsymbol{\alpha})}=\bigl{\{}\mathcal{G}_{p}^{(\boldsymbol{\alpha})}\bigr{\}}_{p=1}^{n_{\text{parts}}},\qquad\mathcal{G}_{p}^{(\boldsymbol{\alpha})}=\bigl{\{}(\boldsymbol{\mu}_{i}^{(\boldsymbol{\alpha})},\boldsymbol{\Sigma}_{i}^{(\boldsymbol{\alpha})},\mathbf{c}_{i},\alpha_{i})\bigr{\}}_{i\in\mathcal{G}_{p}},$ (12)
where $\boldsymbol{\alpha}\in\mathbb{A}$ indexes the interaction state ($|\mathbb{A}|=K$) and $\mathcal{G}_{p}$ denotes the set of Gaussian indices belonging to part $p$. Each state-specific field $\mathcal{G}^{(\boldsymbol{\alpha})}$ encodes a coherent multi-part configuration, while the union over all $\boldsymbol{\alpha}$ compactly describes the full articulated motion range of the object.
@@ 3.5 Multi-Task Modeling for Articulated Objects
Given the part-aware Gaussian field $\mathcal{G}^{(\boldsymbol{\alpha})}=\{\mathcal{G}_{p}^{(\boldsymbol{\alpha})}\}_{p=1}^{n_{\text{parts}}}$, we extract three part-level outputs. First, every subset $\mathcal{G}_{p}^{(\boldsymbol{\alpha})}$ is rendered in depth and color images in all calibrated views; marching cubes applied to the resulting volume produce the mesh $\mathcal{M}_{p}\in\mathbb{R}^{|V_{p}|\times 3}$. Second, to classify the type of joint part $p$, we align two interaction states with the Kabsch algorithm [41] and inspect the residual displacement: a low-rank residual indicates a revolute joint, while a full-rank residual indicates a prismatic joint. Third, the trajectories of the Gaussian centroids belonging to part $p$ are fitted with a minimal motion model. A revolute joint is parameterised by a pivot point $\mathbf{p}\in\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\in\mathbb{R}^{4}$ with $\lVert\mathbf{q}\rVert=1$. A prismatic joint is parameterised by a unit slide axis $\mathbf{a}\in\mathbb{R}^{3}$ with $\lVert\mathbf{a}\rVert=1$ and a translation distance $d\in\mathbb{R}$. More details are provided in Sec. A of the Appendix.
@@ 5 Experiments
Table 1: Results on the Partnet-Mobility Dataset. A superscript denotes that the method failed to obtain a valid result for at least one object in the corresponding category.
Metric Method Manual Synthetic Objects Mean
Box Door Eyeglasses Faucet Oven Fridge Storage Table
Axis $\downarrow$ Ang 0 PARIS [20] 36.27 24.83 64.35 13.75 54.88 69.12 27.05 15.60 38.23
ArticulatedGS [21] 0.98 1.43 2.25 5.14 4.13 14.14 23.24 1.35 6.58
DTArt [23] 20.65 0.38 18.18 1.06 3.37 40.99 0.46 0.40 10.69
ArtGS [24] 0.01 0.35 0.06 8.05 2.33 0.13 6.05 1.98 2.37
Ours 0.49 0.31 0.06 0.88 0.34 0.09 0.21 0.25 0.33
[…]
ArticulatedGS [21] 5.12 0.77 2.12 4.12 5.25 2.13 0.12 0.52 2.52
DTArt [23] 42.60 0.75 1.62 1.44 14.21 2.17 0.22 18.12 10.14
ArtGS [24] 1.36 0.68 0.08 17.02 9.24 1.41 0.01 7.66 4.68
Ours 0.28 0.49 1.22 0.58 0.12 0.73 0.04 0.28 0.47
[…]
ArticulatedGS [21] 0.97 0.26 2.11 0.24 3.46 0.26 0.74 0.79 1.10
DTArt [23] 7.09 0.01 4.96 0.02 14.89 11.35 0.07 0.03 4.80
ArtGS [24] 0.06 0.03 0.06 0.01 1.9 0.02 1.28 39.14 5.31
Ours 0.06 0.02 0.02 0.04 0.27 0.16 0.03 0.01 0.08
[…]
ArticulatedGS [21] 0.24 0.62 0.24 1.24 2.15 0.15 0.24 0.65 0.69
DTArt [23] 3.96 0.21 11.75 0.02 32.87 0.15 0.01 0.01 6.12
ArtGS [24] 0.51 0.03 0.04 0.05 6.24 0.04 0.01 5.98 1.61
Ours 0.00 0.02 0.09 0.04 0.08 0.03 0.00 0.01 0.03
[…]
ArticulatedGS [21] 4.24 2.65 34.24 2.46 6.35 3.35 2.33 0.36 7.00
DTArt [23] 42.85 0.35 13.91 0.07 70.02 36.73 0.07 0.55 20.57
ArtGS [24] 2.78 0.05 0.07 6.80 1.67 0.03 6.69 8.97 3.38
Ours 2.09 0.05 0.02 0.95 0.74 0.10 0.43 0.32 0.59
[…]
ArticulatedGS [21] 20.04 7.24 0.63 9.24 2.58 19.87 0.90 0.23 7.59
DTArt [23] 63.48 89.31 104.92 1.44 65.79 4.91 0.05 0.21 41.26
ArtGS [24] 15.68 0.05 0.08 16.89 3.87 0.03 0.16 24.09 7.60
Ours 0.87 0.05 0.02 0.50 0.34 0.02 0.21 0.06 0.26
CD-s $\downarrow$ PARIS [20] 6.93 17.90 23.35 9.36 13.45 18.96 10.44 3.45 12.98
ArticulatedGS [21] 18.17 167.53 15.14 0.97 29.43 23.80 16.18 5.92 34.64
DTArt [23] 2.45 20.60 49.28 0.49 2.60 4.06 1.27 8.65 11.17
ArtGS [24] 3.48 0.35 0.13 1.05 5.44 9.69 5.10 9.04 4.29
Ours 1.97 0.80 0.11 0.64 1.66 4.15 2.78 4.27 2.05
CD-m 0 $\downarrow$ PARIS [20] 82.05 42.22 86.37 34.79 107.25 212.08 182.85 16.58 95.52
ArticulatedGS [21] 9.65 0.84 26.88 35.63 1.91 11.24 7.32 7.24 12.59
DTArt [23] 363.75 0.22 47.24 0.22 157.26 0.63 0.23 46.30 76.98
ArtGS [24] 2.47 0.44 0.16 0.38 8.89 0.86 3.46 143.78 20.06
Ours 2.44 0.15 0.09 0.71 1.79 0.82 0.03 5.40 1.43
CD-m 1 $\downarrow$ PARIS [20] 128.58 32.08 98.54 48.86 301.18 76.82 125.06 76.18 110.91
ArticulatedGS [21] 8.78 0.44 13.47 0.27 10.92 1.15 179.27 12.82 28.39
DTArt [23] 198.70 1.17 1133.05 0.22 140.52 0.36 0.27 0.22 184.31
ArtGS [24] 3.35 0.42 0.17 0.46 8.47 0.86 2.25 1145.91 145.23
Ours 1.84 0.43 0.13 0.74 2.15 0.27 0.11 0.76 0.80
CD-w $\downarrow$ PARIS [20] 6.30 12.59 17.24 9.98 10.93 18.89 8.86 2.81 10.95
ArticulatedGS [21] 13.75 12.59 19.03 2.78 23.22 18.89 50.59 5.81 18.33
DTArt [23] 1.08 1.40 0.15 0.50 1.73 0.98 10.17 1.39 2.18
ArtGS [24] 5.91 1.56 0.17 0.58 7.35 16.82 8.66 12.73 6.72
Ours 4.07 1.24 0.13 0.47 1.48 10.16 5.48 1.14 3.02

### v2

@@ 3.3 Coarse Part-Level Segmentation of Gaussian Primitives
Motion-driven separation of static and dynamic Gaussians. Without object masks or semantic priors, we detect dynamic primitives by computing the maximum Euclidean displacement of each Gaussian center across the $K$ interaction states. Let $\bm{\mu}_{i}^{(k)}\in\mathbb{R}^{3}$ be the center of Gaussian primitives $g_{i}$ in state $k$. We define its maximal displacement as
$d_{i}=\max_{j,k\in\{1,\dots,K\}}\bigl\lVert\bm{\mu}_{i}^{(j)}-\bm{\mu}_{i}^{(k)}\bigr\rVert_{2},\quad\hat{d}_{i}=d_{i}/\!\max_{r}d_{r}\in[0,1].$ (8)
The Gaussians with $\hat{d}_{i}\geq\tau_{\text{mot}}$ form the dynamic set $\mathcal{G}_{\text{dyn}}$, and the remainder $\mathcal{G}_{\text{stat}}$ is static, where the threshold $\tau_{\text{mot}}\in[0,1]$ specifies the fraction of the maximum displacement of the entire scene that qualifies as motion.
Estimating the number of motion parts via a VLM.  To infer how many rigid parts appear in $\mathcal{G}_{\text{dyn}}$, we sample $M$ image pairs $(I_{k_{i}}^{(v_{p})},I_{k_{j}}^{(v_{q})})$ with different states $k_{i}\neq k_{j}$. A visual language model (VLM) (Dai et al., 2023) is queried with the fixed prompt:
@@
Trajectory-based clustering of dynamic Gaussians. For each $g_{i}\!\in\!\mathcal{G}_{\text{dyn}}$ we assemble a motion descriptor
$\mathbf{f}_{i}=\bigl[\hat{\Delta}\bm{\mu}_{i}^{(1)},\,\lVert\Delta\bm{\mu}_{i}^{(1)}\rVert_{2},\,\dots,\,\hat{\Delta}\bm{\mu}_{i}^{(K-1)},\,\lVert\Delta\bm{\mu}_{i}^{(K-1)}\rVert_{2}\bigr]\in\mathbb{R}^{4(K-1)},$ (9)
@@
$\Delta\bm{\mu}_{i}^{(k)}=\bm{\mu}_{i}^{(k+1)}-\bm{\mu}_{i}^{(k)},\qquad\hat{\Delta}\bm{\mu}_{i}^{(k)}=\frac{\Delta\bm{\mu}_{i}^{(k)}}{\lVert\Delta\bm{\mu}_{i}^{(k)}\rVert_{2}+\varepsilon}.$ (10)
Unit directions encode orientation and lengths keep relative travel, so the descriptor is translation/rotation aware yet scale balanced. After normalization of $\ell_{2}$, we cluster the descriptors with K-means (Arthur and Vassilvitskii, 2006) using $K=n_{\text{parts}}$. On the unit sphere, Euclidean distance equals cosine distance, therefore Gaussians that share one rigid motion, even at different amplitudes, are grouped together.
Clusters containing fewer than $2\%$ of dynamic Gaussians are regarded noise and merged into the nearest larger cluster. Each dynamic primitive then receives a part label $c_{i}\in\{1,\dots,n_{\text{parts}}\}$, while static ones are labeled 0. These coarse labels initialize the fine-grained stage that follows.
@@ 3.4 Refined Part-Level Segmentation of Gaussian Primitives
Visibility-aware prompt generation for SAM. To sharpen the coarse part labels in Sec. 3.3 we first render each part through a visibility filter and then draw sparse point prompts for the Segment Anything Model (SAM) (Kirillov et al., 2023).
  • Per-Gaussian contribution at a pixel. For a pixel $\mathbf{x}$ in view $v$, the contribution weight $w_{i}^{(v)}(\mathbf{x})$ from Gaussian $g_{i}=(\bm{\mu}_{i},\bm{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$ is obtained by first evaluating its screen-space density $\rho_{i}^{(v)}(\mathbf{x})$ (Eq. 2) and then applying the front-to-back compositing rule (Eq. 3), which together map every three-dimensional primitive to a continuous weight field whose value at $\mathbf{x}$ equals $w_{i}^{(v)}(\mathbf{x})$.
@@
    where $\tau_{\text{vis}}$ is a visibility threshold, tuned on one or two representative instances, and $\varepsilon$ avoids division by zero. The winner margin test keeps only pixels in which part $p$ clearly dominates, thus reducing the ambiguity at overlaps.
@@
Boundary-aware Gaussian splitting. For each Gaussian $g_{i}$ we find its nearest view $v_{\text{near}}$ by ray sampling and project the ellipsoid to a two-dimensional ellipse. If the ellipse major axis extends beyond the mask of the part $\mathcal{M}_{p}^{(v_{\text{near}})}$, the Gaussian becomes a boundary candidate.
We follow SAGD (Hu et al., 2024) to split each candidate, as shown in the bottom part of Fig. 2. The fraction of the major axis that lies inside the mask is measured in image space and the same ratio is applied to the center $\bm{\mu}_{i}$ and the scale vector $\mathbf{s}_{i}$, resulting in two children $g_{i}^{(\text{part})}$ and $g_{i}^{(\text{bg})}$. The part child retains the in-mask portion and the background child captures the overflow. The background child $g_{i}^{(\text{bg})}$ is re-evaluated in its nearest view; if its ellipse still crosses another part mask the split is applied recursively until every descendant lies inside exactly one mask.
Part-aware Gaussian field via local refinement. Finally, parameters of unsplit Gaussians remain fixed, whereas all new children are locally fine-tuned for a few iterations on $\mathcal{L}_{\text{total}}$ (updating $\mathbf{q}$ and $\alpha$ only) to restore the photo consistency. The recursive split-and-tune process produces sharply aligned, mask-consistent part boundaries while leaving the converged Gaussian field undisturbed.
The entire procedure yields a refined, part-aware Gaussian field with interaction state
$\mathcal{G}^{(\bm{\alpha})}=\bigl\{\mathcal{G}_{p}^{(\bm{\alpha})}\bigr\}_{p=1}^{n_{\text{parts}}},\qquad\mathcal{G}_{p}^{(\bm{\alpha})}=\bigl\{(\bm{\mu}_{i}^{(\bm{\alpha})},\bm{\Sigma}_{i}^{(\bm{\alpha})},\mathbf{c}_{i},\alpha_{i})\bigr\}_{i\in\mathcal{G}_{p}},$ (12)
where $\bm{\alpha}\in\mathbb{A}$ indexes the interaction state ($|\mathbb{A}|=K$) and $\mathcal{G}_{p}$ denotes the set of Gaussian indices belonging to part $p$. Each state-specific field $\mathcal{G}^{(\bm{\alpha})}$ encodes a coherent multi-part configuration, while the union over all $\bm{\alpha}$ compactly describes the full articulated motion range of the object.
@@ 3.5 Multi-Task Modeling for Articulated Objects
Given the part-aware Gaussian field $\mathcal{G}^{(\bm{\alpha})}=\{\mathcal{G}_{p}^{(\bm{\alpha})}\}_{p=1}^{n_{\text{parts}}}$, we extract three part-level outputs. First, every subset $\mathcal{G}_{p}^{(\bm{\alpha})}$ is rendered in depth and color images in all calibrated views; marching cubes applied to the resulting volume produce the mesh $\mathcal{M}_{p}\in\mathbb{R}^{|V_{p}|\times 3}$. Second, to classify the type of joint part $p$, we align two interaction states with the Kabsch algorithm (Lawrence et al., 2019) and inspect the residual displacement: a low-rank residual indicates a revolute joint, while a full-rank residual indicates a prismatic joint. Third, the trajectories of the Gaussian centroids belonging to part $p$ are fitted with a minimal motion model. A revolute joint is parameterized by a pivot point $\mathbf{p}\in\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\in\mathbb{R}^{4}$ with $\lVert\mathbf{q}\rVert=1$. A prismatic joint is parameterized by a unit slide axis $\mathbf{a}\in\mathbb{R}^{3}$ with $\lVert\mathbf{a}\rVert=1$ and a translation distance $d\in\mathbb{R}$. More details are provided in Sec. A of the Appendix.
@@ 4 The RS-Art Dataset
Table 1: Results on the PartNet-Mobility Dataset. A superscript denotes that the method failed to obtain a valid result for at least one object in the corresponding category. Note that DTArt requires additional depth input, while other methods rely on RGB inputs.
Metric Method Synthetic Objects Mean
Box Door Eyeglasses Faucet Oven Fridge Storage Table
Axis $\downarrow$ Ang 0 PARIS (Liu et al., 2023) 36.27 24.83 64.35 13.75 54.88 69.12 27.05 15.60 38.23
ArticulatedGS (Guo et al., 2025) 0.98 1.43 2.25 5.14 4.13 14.14 23.24 1.35 6.58
DTArt (Weng et al., 2024) 20.65 0.38 18.18 1.06 3.37 40.99 0.46 0.40 10.69
ArtGS (Liu et al., 2025b) 0.01 0.35 0.06 8.05 2.33 0.13 6.05 1.98 2.37
PD2GS (Ours) 0.49 0.31 0.06 0.88 0.34 0.09 0.21 0.25 0.33
Axis $\downarrow$ Ang 1 PARIS (Liu et al., 2023) 2.89 31.69 3.25 2.65 17.29 49.39 18.00 48.61 21.72
ArticulatedGS (Guo et al., 2025) 5.12 0.77 2.12 4.12 5.25 2.13 0.12 0.52 2.52
DTArt (Weng et al., 2024) 42.60 0.75 1.62 1.44 14.21 2.17 0.22 18.12 10.14
ArtGS (Liu et al., 2025b) 1.36 0.68 0.08 17.02 9.24 1.41 0.01 7.66 4.68
PD2GS (Ours) 0.28 0.49 1.22 0.58 0.12 0.73 0.04 0.28 0.47
Axis $\downarrow$ Pos 0 PARIS (Liu et al., 2023) 0.14 0.50 1.12 0.99 0.29 0.13 0.10 0.29 0.45
ArticulatedGS (Guo et al., 2025) 0.97 0.26 2.11 0.24 3.46 0.26 0.74 0.79 1.10
DTArt (Weng et al., 2024) 7.09 0.01 4.96 0.02 14.89 11.35 0.07 0.03 4.80
ArtGS (Liu et al., 2025b) 0.06 0.03 0.06 0.01 1.9 0.02 1.28 39.14 5.31
PD2GS (Ours) 0.06 0.02 0.02 0.04 0.27 0.16 0.03 0.01 0.08
Axis $\downarrow$ Pos 1 PARIS (Liu et al., 2023) 0.33 0.40 0.77 0.44 0.11 0.18 0.20 0.02 0.31
ArticulatedGS (Guo et al., 2025) 0.24 0.62 0.24 1.24 2.15 0.15 0.24 0.65 0.69
DTArt (Weng et al., 2024) 3.96 0.21 11.75 0.02 32.87 0.15 0.01 0.01 6.12
ArtGS (Liu et al., 2025b) 0.51 0.03 0.04 0.05 6.24 0.04 0.01 5.98 1.61
PD2GS (Ours) 0.00 0.02 0.09 0.04 0.08 0.03 0.00 0.01 0.03
Part $\downarrow$ Motion 0 PARIS (Liu et al., 2023) 93.80 151.50 97.24 65.79 59.28 99.33 45.24 16.38 78.57
ArticulatedGS (Guo et al., 2025) 4.24 2.65 34.24 2.46 6.35 3.35 2.33 0.36 7.00
DTArt (Weng et al., 2024) 42.85 0.35 13.91 0.07 70.02 36.73 0.07 0.55 20.57
ArtGS (Liu et al., 2025b) 2.78 0.05 0.07 6.80 1.67 0.03 6.69 8.97 3.38
PD2GS (Ours) 2.09 0.05 0.02 0.95 0.74 0.10 0.43 0.32 0.59
Part $\downarrow$ Motion 1 PARIS (Liu et al., 2023) 47.89 33.11 87.44 19.35 28.29 150.42 57.09 81.32 63.11
ArticulatedGS (Guo et al., 2025) 20.04 7.24 0.63 9.24 2.58 19.87 0.90 0.23 7.59
DTArt (Weng et al., 2024) 63.48 89.31 104.92 1.44 65.79 4.91 0.05 0.21 41.26
ArtGS (Liu et al., 2025b) 15.68 0.05 0.08 16.89 3.87 0.03 0.16 24.09 7.60
PD2GS (Ours) 0.87 0.05 0.02 0.50 0.34 0.02 0.21 0.06 0.26
CD-s $\downarrow$ PARIS (Liu et al., 2023) 6.93 17.90 23.35 9.36 13.45 18.96 10.44 3.45 12.98
ArticulatedGS (Guo et al., 2025) 18.17 167.53 15.14 0.97 29.43 23.80 16.18 5.92 34.64
DTArt (Weng et al., 2024) 2.45 20.60 49.28 0.49 2.60 4.06 1.27 8.65 11.17
ArtGS (Liu et al., 2025b) 3.48 0.35 0.13 1.05 5.44 9.69 5.10 9.04 4.29
PD2GS (Ours) 1.97 0.80 0.11 0.64 1.66 4.15 2.78 4.27 2.05
CD-m 0 $\downarrow$ PARIS (Liu et al., 2023) 82.05 42.22 86.37 34.79 107.25 212.08 182.85 16.58 95.52
ArticulatedGS (Guo et al., 2025) 9.65 0.84 26.88 35.63 1.91 11.24 7.32 7.24 12.59
DTArt (Weng et al., 2024) 363.75 0.22 47.24 0.22 157.26 0.63 0.23 46.30 76.98
ArtGS (Liu et al., 2025b) 2.47 0.44 0.16 0.38 8.89 0.86 3.46 143.78 20.06
PD2GS (Ours) 2.44 0.15 0.09 0.71 1.79 0.82 0.03 5.40 1.43
CD-m 1 $\downarrow$ PARIS (Liu et al., 2023) 128.58 32.08 98.54 48.86 301.18 76.82 125.06 76.18 110.91
ArticulatedGS (Guo et al., 2025) 8.78 0.44 13.47 0.27 10.92 1.15 179.27 12.82 28.39
DTArt (Weng et al., 2024) 198.70 1.17 1133.05 0.22 140.52 0.36 0.27 0.22 184.31
ArtGS (Liu et al., 2025b) 3.35 0.42 0.17 0.46 8.47 0.86 2.25 1145.91 145.23
PD2GS (Ours) 1.84 0.43 0.13 0.74 2.15 0.27 0.11 0.76 0.80
CD-w $\downarrow$ PARIS (Liu et al., 2023) 6.30 12.59 17.24 9.98 10.93 18.89 8.86 2.81 10.95
ArticulatedGS (Guo et al., 2025) 13.75 12.59 19.03 2.78 23.22 18.89 50.59 5.81 18.33
DTArt (Weng et al., 2024) 1.08 1.40 0.15 0.50 1.73 0.98 10.17 1.39 2.18
ArtGS (Liu et al., 2025b) 5.91 1.56 0.17 0.58 7.35 16.82 8.66 12.73 6.72
PD2GS (Ours) 4.07 1.24 0.13 0.47 1.48 10.16 5.48 1.14 3.02